A new flag is to be designed with 6 vertical strip using some or all of the colour yellow,green,blue and red . Then the no. Of ways this can be done such that no two adjacent strips have the same colour is

Grade:11

4 Answers

Gosula Madhukar
8 years ago
there are 972 ways.
i think iam right
if u don’t agree please send the answer
thank you
bye.
if u have any further doubts plz post them
 
Ankit Jaiswal
8 years ago
u did not provide any method or procedure about the solution kindly send the solution too
I ALSO HAVE THIS DOUBT 
PLZ ANYONE HELP US
Gundeep Singh
8 years ago
The ans. Is 972 ways. In general for a flag of n colored strips number of ways = 4×3^(n-1). The method is long or can`t be typed...and i think there is no option for sending image files in answers. Just try to alalyse the question by first taking only 1 stip in a flag , then 2strips and so on....
Gosula Madhukar
8 years ago
LET A,B,C,D,E,F BE THE POSITIONS OF THE STRIPS.
THEN POSITION A HAS 4 POSSIBILITIES,WHERE AS THE REST HAS THREE POSSIBILITIES EACH.
THEREFORE THE NO. OF WAYS(BY THE RULE OF ‘AND’)=4*3*3*3*3*3=972.